# Violation of Bell inequalities in an analogue black hole

### v1

@@ I Introduction
The domain of analog gravity aims at providing laboratory models for gaining insight on general relativity phenomena that cannot be directly observed in the usual gravitational context or for which there exists no complete theoretical framework. Two such phenomena are black hole superradiance [1, 2, 3] and Hawking radiation [4, 5]. In that context, the most successful analogous experimental platforms have been surface gravity [6, 7, 8, 9, 10] and acoustic [11] waves, nonlinear optical systems [12, 13], cavity polaritons [14, 15, 16] and Bose-Einstein condendates (BEC) of atomic vapors [17, 18, 19]. Because of their low temperature and coherence properties BECs appear particularly well suited for demonstrating quantum features, for instance for deciding whether the phenomenon of analogous Hawking radiation is mostly stimulated or does actually correspond to spontaneous emission of pairs of entangled (quasi-)particles as in Hawking’s original scenario. We do not directly address this point in the present work (see, e.g., Ref. [20] for a recent discussion) but rather take the theoretical analysis a little further by asking: which general insight can we reach by studying quantum correlations of the Hawking signal emitted by an analogue black hole? A natural approach for such an investigation is a test of nonlocality via violation of Bell inequality. The epistemological query of refutation of local hidden variable theories has already received an unambiguous answer in many contexts (see, e.g., Ref. [21] and references therein) and important progresses have also been achieved in the field of BEC matter waves [22, 23, 24, 25] we consider here. Such a test is nonetheless a nontrivial extension of the scope of analogue gravity and would constitute a primer for continuous variables entanglement in a matter wave environment (see also Refs. [26, 27, 28] for related proposals). In view of future experimental studies it is relevant to quantitatively evaluate in realistic configurations to what extent Bell inequalities can be violated in BEC analogues. This is a natural question to ask, all the more so as we argue in the following that in some (exotic) limits the analog black hole we consider exactly realizes a Einstein-Podolsky-Rosen (EPR) pair. Furthermore, as will be shown, the specifics of the system provide a natural testing ground for genuine tripartite nonlocality. Our theoretical investigation of the matter reveals an unexpected generic feature of black hole analogues: in the long wave-length limit, the state of the system realizes an infinite sum of degenerate Greenberger-Horne-Zeilinger (GHZ) states. Interestingly, thanks to the continuous nature of its degrees of freedom, and despite the clear GHZ nature of its long wavelength modes, the analog system remains entangled after partial tracing.
@@ II The model
where $\Phi(x)$ describes the background flow while $\hat{\psi}$ accounts for small quantum fluctuations. Whereas the decomposition (2) is legitimate in 3D, it has a finite range of validity in the 1D configurations we consider; however, its conditions of applicability are commonly met in standard experimental situations, see, e.g., the discussion in [30]. In decomposition (2) the classical field $\Phi$ is a stationary solution of the Gross-Pitaevskii equation, see Eq. (35). The associated flow realizes an analogue black hole horizon if it is subsonic in the asymptotic upstream region and supersonic in the asymptotic downstream region. Several such flows have been considered in the past. We will here focus on the “waterfall configuration” which is close to the experimental realization of Refs. [31, 18]. In order to assess the generality of our results, we present in Appendix I the results obtained for two other configurations, we denote as “delta peak” and “flat profile”. All these configurations are described in Appendix A.
@@ II.1 Propagation channels and quantum modes
The decomposition (2) is meaningful in a regime of small quantum fluctuations where the operator $\hat{\psi}$ can be treated within a Bogoliubov approach. In this case $\hat{\psi}$ is naturally expanded along the asymptotic ingoing and outgoing channels of the flow. The dispersion relation of elementary excitations in the asymptotic upstream subsonic and downstream supersonic regions ($x\to-\infty$ and +$\infty$+ ∞, respectively) is of Bogoliubov type, with a Doppler shift accounting for the finite velocity of the background:
@@ II.2 Density matrix and measure of bipartite entanglement
where $n_{\rm\scriptscriptstyle th}(\omega)$ is the thermal Bose occupation distribution at temperature $T$ and energy $\hbar\omega$, whereas $\omega_{{\scriptscriptstyle\rm B},\alpha}(q_{j|\rm in})$ is the Bogoliubov dispersion relation (4), with $\alpha=u$ if $j=0$ and $\alpha=d$ if $j=1$ or 2. The functions $q_{j|\rm in}(\omega)$ appearing in expression (14) are pictorially defined in Fig. 1. For instance $q_{2|\rm in}(\omega)$ is the function that, to a given angular frequency $\omega\in[0,\Omega]$, associates a wave-vector along the $2|\rm in$ dispersion branch111 $q_{2|\rm in}(\omega)\in[q^{*},q_{0}]$ with $q^{*}=q_{2|\rm in}(\Omega)$ and $q_{0}=q_{2|\rm in}(0)$, see Fig. 1.. We loosely refer to the cases where $\bar{n}_{j}$ ($j\in\{0,1,2\}$) is equal to the right-hand side (r.h.s.) of (14) as “finite temperature” situations. In the simplest configuration, denoted as “zero temperature” the system is in the pure state $|0\rangle^{\rm in}$ and the $\bar{n}_{j}$’s are all equal to zero. We will not specify the values of the $\bar{n}_{j}$’s in the following, so that the formulas we give are generally valid, even in situations where the occupation numbers should not be given by formulas of the type of Eq. (14). However, for illustrative purposes, all the figures of the paper are plotted in specific temperature cases.
Several observables have been proposed to theoretically evaluate the bipartite entanglement in the context of analogue gravity, such as the Cauchy-Schwarz criterion [38, 39, 40, 41, 42, 43], the generalized Peres-Horodecki parameter [44, 45], the logarithmic negativity [46, 47, 48, 49], the entanglement entropy [50, 49], the entanglement of formation [51], the Gaussian contangle [20]. In the present work, for reasons exposed in Appendix B, we chose to evaluate the bipartite entanglement between modes $i$ and $j$ by means of a quantity $\Lambda^{(i|j)}(\omega)\in]-\infty,1]$ which is a monotonous measure of entanglement that has been denoted as the “PPT measure” in Ref. [20]. States $i$ and $j$ are separable when $\Lambda^{(i|j)}<0$, which is always the case when $(i|j)=(0|1)$: the companion and the Hawking modes are not entangled. The two other couples of modes, $(0|2)$ and $(1|2)$, are always entangled at $T=0$ for all $\omega$. Their entanglement decreases with increasing temperature by an amount specified by the PPT measure which reads explicitly ($i=0$ or 1) [20]:
@@ III Bipartite nonlocality
Most Bell-like inequalities proposed in the context of continuous variables hinge on a discretization process [54, 55, 56, 57, 58, 59]. Indeed, since the set of outcomes for a given observable is typically unbounded in continuous variable systems, it seems a priori difficult to derive an upper bound of the expectation value of a Bell-type observable. It is not impossible, though: relying on the Fine–Abramsky–Brandenburger theorem [60, 61], the authors of Refs. [62, 63] were able to derive continuous Bell inequalities for continuous and unbounded observables. However, to our knowledge, no practical (theoretical or experimental) use of such fully continuous approaches has been successfully implement so far. A method to circumvent this difficulty is based on discretization and consists of using a dichotomic binning of the outcome results. In this case, by using observables which rely on continuous measurements but can only take a finite number of outcomes, one can construct a Bell inequality similar to those derived for discrete variables. A well-known example of such observables are the so-called pseudo-spin operators: they live in an infinite-dimensional Hilbert space but the outcome of their measurement is either $-1$ or $+1$.
@@
$\hat{\Pi}^{(j)}_{x}(\omega)=\int_{0}^{+\infty}{\rm d}q\,\Big{(}|q\rangle_{j}\,{}_{j}\!\langle q|-|-q\rangle_{j}\,{}_{j}\langle-q|\Big{)},$ (16a)
$\hat{\Pi}^{(j)}_{y}(\omega)=i\int_{0}^{+\infty}{\rm d}q\,\Big{(}|q\rangle_{j}\,{}_{j}\langle-q|-|-q\rangle_{j}\,{}_{j}\langle q|\Big{)},$ (16b)
@@
^$\begin{split}\hat{\mathscr{B}}{}^{(i|j)}(\omega)=&(\boldsymbol{a}+\boldsymbol{a}^{\prime})\cdot\hat{\boldsymbol{\Pi}}{}^{(i)}\otimes\boldsymbol{b}\cdot\hat{\boldsymbol{\Pi}}{}^{(j)}\\ +&(\boldsymbol{a}-\boldsymbol{a}^{\prime})\cdot\hat{\boldsymbol{\Pi}}{}^{(i)}\otimes\boldsymbol{b}^{\prime}\cdot\hat{\boldsymbol{\Pi}}{}^{(j)},\end{split}$ (18)
where $\boldsymbol{a}$, $\boldsymbol{a}^{\prime}$, $\boldsymbol{b}$ and $\boldsymbol{b}^{\prime}$ are unit vectors. Given a unit vector $\boldsymbol{n}$ it is easily checked that $(\boldsymbol{n}\cdot\hat{\boldsymbol{\Pi}}{}^{(j)})^{2}=\mathbb{1}$, meaning that the Hermitian operator $\boldsymbol{n}\cdot\hat{\boldsymbol{\Pi}}{}^{(j)}$ has eigenvalues ±$\pm 1$. It then follows from direct inspection that, if one believes in local realism, one should expect that a measure of the operator ^$\hat{\mathscr{B}}{}^{(i|j)}$ yields a result ±$\pm 2$ [70]. The tenant of local realism is thus violated when ^$\langle\hat{\mathscr{B}}{}^{(i|j)}\rangle>2$, whereas Cirel’son bound [71] imposes ^$\langle\hat{\mathscr{B}}{}^{(i|j)}\rangle\leq 2\sqrt{2}$. Since the modes 0 and 1 are not entangled the quantity ^$\langle\hat{\mathscr{B}}{}^{(0|1)}\rangle$ is always lower than 2 and its computation is of no interest to us. For attempting to violate as much as possible Bell inequality one should consider the modes 0 and 2 (or 1 and 2) and look for an arrangement of the four measurement directions $\boldsymbol{a}$, $\boldsymbol{a}^{\prime}$, $\boldsymbol{b}$ and $\boldsymbol{b}^{\prime}$ which maximizes ^$\langle\hat{\mathscr{B}}{}^{(i|2)}\rangle$ ($i=0$ or 1). This procedure is explained in Appendix F and makes it possible to analytically compute the quantity
@@
Figure 3: Plot of $B^{(i|2)}$ (solid lines) and $\Lambda^{(i|2)}$ (dashed lines) as functions of $\omega$ for the waterfall configuration with $m_{d}=2.9$ and different temperatures. We only consider the range of frequencies $\omega<\Omega$ for which the vacuum of the outgoing and ingoing modes do not coincide. The value of the temperature is indicated in units of $g\,n_{u}=mc_{u}^{2}$. Upper plot: $i=0$, lower plot $i=1$. Non separability of modes $i$ and $2$ is achieved when $\Lambda^{(i|2)}>0$. Bell inequality is violated when $B^{(i|2)}>2$.
@@
Fig. 3 indicates that, in the waterfall configuration we consider (with $m_{d}=2.9$), the entanglement is lower and the violation of Bell inequality less significant for the correlations among modes 1 and 2 than for the correlations among modes 0 and 2. However, this is not always the case. This point is illustrated in Fig. 4 which displays, for all waterfall configurations, the zero temperature values of ${\rm max}_{\omega}B^{(i|2)}$ and ${\rm max}_{\omega}\Lambda^{(i|2)}$ (for $i=0$ and 1) as functions of the upper Mach number $m_{u}$ (the maximisation is performed at fixed $m_{u}$, for $\omega\in[0,\Omega]$). All the possible waterfall configurations are considered since $m_{u}$ spans the whole interval $[0,1]$. In this figure we aim at evaluating the largest amount of entanglement and nonlocality reached in each configuration. This is the reason why we plot the maximum values taken by the quantities $\Lambda^{(i|2)}$ and $B^{(i|2)}$ over the energy interval $[0,\Omega]$ [where $\Omega$ is defined by Relation (7)] since this is only for energies in this interval that spontaneous emission of quasi-particles occurs.
Figure 4: Zero temperature value of the CHSH parameter and of the PPT measure characterizing non separability of modes $i$ ($=0$ and 1) and 2 for the waterfall configuration. The maximal value reached by these quantities over the interval $\omega\in[0,\Omega]$ is plotted as a function of the upper Mach number $m_{u}$ which (as explained in Appendix A) characterizes a given configuration. The values of $B^{(1|2)}$ for $m_{u}\leq 0.01$ are not indicated because of lack of numerical precision. The upper bounds of $B^{(i|2)}$ and $\Lambda^{(i|2)}$ ($\sqrt{8}$ and 1, respectively) are indicated with filled dots.
@@
Fig. 4 shows that for values of $m_{u}$ larger than 0.6, the entanglement is mainly concentrated between modes 2 and 0, i.e., between the Hawking quantum and the partner. This is indicated by the fact that both the PPT measure and the CHSH parameter significantly point to nonseparability and nonlocality between these two modes. For $m_{u}\leq 0.2$ instead, the figure shows that entanglement is concentrated between the partner and the companion (modes 2 and 1, respectively).
@@
A plot similar to the one of Fig. 4, but where the quantities are evaluated at finite temperature, enables to evaluate the resilience of entanglement and non-separability to an increase of temperature. This check is performed in Fig. 5 which shows that, whereas at $T=0.2\,gn_{u}$ the PPT measure is not dramatically affected, the CHSH parameters $B^{(0|2)}$ and $B^{(1|2)}$ no longer show evidences of violation of Bell inequality, except in the $(1|2)$ sector for waterfall configurations with $m_{u}\lesssim 0.15$ and, in a lesser extent, in the $(0|2)$ sector for $m_{u}\gtrsim 0.85$.
@@
where ^$\hat{\mathscr{B}}^{\prime}{}^{(j|k)}(\omega)$ is the same as ^$\hat{\mathscr{B}}{}^{(j|k)}(\omega)$ defined in (18) with the primes reversed, and $\boldsymbol{c}$ and $\boldsymbol{c}^{\prime}$ are normalized vectors (as well as $\boldsymbol{a}$, $\boldsymbol{a}^{\prime}$, $\boldsymbol{b}$, and $\boldsymbol{b}^{\prime}$ involved in the definition of ^$\hat{\mathscr{B}}{}^{(j|k)}$). Expanding expression (20) shows that $\hat{\mathscr{S}}{}^{(i|j|k)}$ is invariant upon a permutation of its indices, provided the names of the unit vectors $(\boldsymbol{a},\boldsymbol{b},\boldsymbol{c})$ and $(\boldsymbol{a}^{\prime},\boldsymbol{b}^{\prime},\boldsymbol{c}^{\prime})$ undergo the same permutation. In the following we arbitrarily chose the order $(i,j,k)=(0,1,2)$.
Similarly to what occurs for the two-mode operator, the principle of local realism, if correct, should predict that a measure of the operator $\hat{\mathscr{S}}{}^{(0|1|2)}$ yields results ±$\pm 2$. The system violates this principle when the average of the operator (20) is larger than two. This is often referred to as violation of Svetlichny inequality. It is important to note that the observable $\hat{\mathscr{S}}{}^{(0|1|2)}$ is specially designed in order to be sensitive to genuine tripartite nonlocality (see discussions in Refs. [74, 76, 75, 21] and references therein): a tripartite system can involve nonlocal correlations between any of its bipartitions and still not violate the Svetlichny inequality $\langle\hat{\mathscr{S}}{}^{(0|1|2)}\rangle<2$. A simple example of a system which displays two mode nonlocality but does not pass the Svetlichny test is presented in Appendix C [see Eq. (64) and the discussion below].
@@
Figure 7: Same as Fig. 6 for a temperature $T=0.05\,gn_{u}$. At variance with the zero temperature situation the tripartite measure (21) is here always lower or equal to 2.
@@
All the other expectation values of products of three components of the $\hat{\boldsymbol{\Pi}}$ operator are zero. Since all the operators in (22) have only ±$\pm 1$ as eigenvalues, this means that each term must reach its extremal value and the vacuum mode of zero frequency $|0_{\omega=0}\rangle^{\rm in}$ must therefore be an eigenstate of the operators appearing in (22):
@@
Relations (23) define a state which exhibits the GHZ paradox, contradicting local hidden variable theories by means of a single measurement; cf. e.g., the discussion in [77]. In the same line, the above noted fact that our system saturates the upper bound $2\sqrt{2}$ of the Svetlichny parameter is a characteristic also shared by GHZ states. Another resemblance lies in the fact that the GHZ states possess maximal tripartite entanglement, quantified by the residual tangle [78, 79] while similarly, in the domain of temperature and energy where Eqs. (23) hold, the residual contangle [80] which measures genuine tripartite entanglement in our continuous variable system diverges [20]. Despite these clear similarities, there is a significant difference between the state $|0_{\omega=0}\rangle^{\rm in}$ and the GHZ states commonly considered in quantum information theory: they have different behaviors upon partial tracing. As well known, taking the partial trace over one of the three modes of a GHZ state yields an unentangled mixed state. At variance, partial tracing the state $|0_{\omega=0}\rangle^{\rm in}$ over modes 0 or 1 leads to an entangled two-mode mixed state, as discussed in Sec. III (see, e.g., the values $\Lambda^{(0|2)}(\omega=0)$ and $\Lambda^{(1|2)}(\omega=0)$ in Fig. 3). As we now argue, the explanation of this conundrum lies in the fact that, whereas the GHZ states are usually build on qubits, the state we consider is an infinite sum of degenerate GHZ states of a continuous variable system.
To tackle this issue it is convenient to expand the state $|0_{\omega=0}\rangle^{\rm in}$ over the eigenstates of the operators $\hat{\Pi}_{x}^{(j)}$ ($j=0$, 1 or 2). As discussed in Appendix D these eigenstates can be written as $|x^{\pm}_{n}\rangle_{j}$: they are labeled by their eigenvalue (±$\pm 1$) plus another integer index ($n$ in the above expression) associated to the infinite degeneracy of both eigenvalues. Indeed, at variance with what occurs for a regular spin operator, the projection of the pseudo-spin (16) over a given axis (here $\hat{\Pi}_{x}$) has infinitely degenerate eigenvalues, i.e., there exists an infinite number of mutually orthogonal eigenstates with the same eigenvalue ($+1$ or $-1$):
@@
$\begin{split}|0_{\omega=0}\rangle^{\rm in}=\sum_{l,m,n}C_{l,m,n}\Big{(}&|x^{+}_{l},x^{-}_{m},x^{-}_{n}\rangle\,+\\ &|x^{-}_{l},x^{+}_{m},x^{+}_{n}\rangle\Big{)},\end{split}$ (29)

### v2

@@ I Introduction
The domain of analog gravity aims at providing laboratory models for gaining insight on general relativity phenomena that cannot be directly observed in the usual gravitational context or for which there exists no complete theoretical framework. Two such phenomena are black hole superradiance [1, 2, 3] and Hawking radiation [4, 5]. In that context, the most successful analogous experimental platforms have been surface gravity [6, 7, 8, 9, 10] and acoustic [11] waves, nonlinear optical systems [12, 13, 14, 15], cavity polaritons [16, 17, 18] and Bose-Einstein condendates (BEC) of atomic vapors [19, 20, 21]. Because of their low temperature and coherence properties BECs appear particularly well suited for demonstrating quantum features. We will not address here the question of the actual experimental demonstration of quantum entanglement in a BEC analogue black hole (see, e.g., Ref. [22] for a recent discussion) but rather take the theoretical analysis a little further by asking: which general insight can we reach by studying quantum correlations of the Hawking signal emitted by an analogue black hole? A natural approach for such an investigation is a test of nonlocality via violation of Bell inequality. The epistemological query of refutation of local hidden variable theories has already received an unambiguous answer in many contexts (see, e.g., Ref. [23] and references therein) and important progresses have also been achieved in the field of BEC matter waves [24, 25, 26, 27] we consider here. Such a test is nonetheless a nontrivial extension of the scope of analogue gravity and would constitute a primer for continuous variables entanglement in a matter wave environment (see also Refs. [28, 29, 30] for related proposals). In view of future experimental studies it is relevant to quantitatively evaluate in realistic configurations to what extent Bell inequalities can be violated in BEC analogues. This is a natural question to ask, all the more so as we argue in the following that in some (exotic) limits the analog black hole we consider exactly realizes a Einstein-Podolsky-Rosen (EPR) pair. Furthermore, as will be shown, the specifics of the system provide a natural testing ground for genuine tripartite nonlocality. Our theoretical investigation of the matter reveals an unexpected generic feature of black hole analogues: in the long wave-length limit, the state of the system realizes an infinite sum of degenerate Greenberger-Horne-Zeilinger (GHZ) states. Interestingly, thanks to the continuous nature of its degrees of freedom, and despite the clear GHZ nature of its long wavelength modes, the analog system remains entangled after partial tracing.
@@ II The model
where $\Phi(x)$ describes the background flow while $\hat{\psi}$ accounts for small quantum fluctuations. Whereas the decomposition (2) is legitimate in 3D, it has a finite range of validity in the 1D configurations we consider; however, its conditions of applicability are commonly met in standard experimental situations, see, e.g., the discussion in [32] which is summarized in Appendix A. In decomposition (2) the classical field $\Phi$ is a stationary solution of the Gross-Pitaevskii equation, see Eq. (36). The associated flow realizes an analogue black hole horizon if it is subsonic in the asymptotic upstream region and supersonic in the asymptotic downstream region. Several such flows have been considered in the past. We will here focus on the “waterfall configuration” which is close to the experimental realization of Refs. [33, 20]. In order to assess the generality of our results, we present in Appendix I the results obtained for two other configurations, we denote as “delta peak” and “flat profile”. All these configurations are described in Appendix A.
@@ II.1 Propagation channels and quantum modes
The decomposition (2) is meaningful in a regime of small quantum fluctuations where the operator $\hat{\psi}$ can be treated within a Bogoliubov approach. In this case $\hat{\psi}$ is naturally expanded along the asymptotic ingoing and outgoing channels of the flow. The dispersion relation of elementary excitations in the asymptotic upstream subsonic and downstream supersonic regions ($x\to$-$\infty$ and $\infty$, respectively) is of Bogoliubov type, with a Doppler shift accounting for the finite velocity of the background:
@@ II.2 Density matrix and measure of bipartite entanglement
In expression (12) and in all the following, the averages $\langle\cdots\rangle$ are performed over the density matrix of the system. This density matrix is simply $\rho=|0\rangle^{\rm in}\,{}^{\rm in}\langle 0|$ in the ideal case just described. We also consider more realistic situations where some incoherent excitations are present and the system is not in a pure state. A simple manner to account for this situation would be to assume that the system is in a thermal state. This is however impossible because the analog configurations depicted in Appendix A are thermodynamically unstable. A way to circumvent this problem has been proposed in Refs. [39, 40]. It consists in postulating that the system was initially in thermal equilibrium at temperature $T$ with a constant density and velocity ($n_{u}$ and $V_{u}$, respectively) and that the flow has been adiabatically modified by slowly ramping the appropriate external potential, eventually reaching the black hole configuration of interest. This situation, although idealized, it less schematic than the zero excitation regime. It emulates the experimental situation of Refs. [20, 21] if the system is considered to have been initially in equilibrium in the frame attached to the flowing condensate. In this case one has (for $j\in\{0,1,2\}$)
@@
Several observables have been proposed to theoretically evaluate the bipartite entanglement in the context of analogue gravity, such as the Cauchy-Schwarz criterion [41, 42, 43, 44, 45, 46], the generalized Peres-Horodecki parameter [47, 48], the logarithmic negativity [49, 50, 51, 52], the entanglement entropy [53, 52], the entanglement of formation [54], the Gaussian contangle [22]. In the present work, for reasons recalled in Appendix B, we chose as in Ref. [22] to evaluate the bipartite entanglement between modes $i$ and $j$ by means of a quantity $\Lambda^{(i|j)}(\omega)\in]-\infty,1]$ which is a monotonous measure of entanglement that we denote as the “PPT measure”. States $i$ and $j$ are separable when $\Lambda^{(i|j)}<0$, which is always the case when $(i|j)=(0|1)$: the companion and the Hawking modes are not entangled. The two other couples of modes, $(0|2)$ and $(1|2)$, are always entangled at $T=0$ for all $\omega$. Their entanglement decreases with increasing temperature by an amount specified by the PPT measure which reads explicitly ($i=0$ or 1) [22]:
@@ III Bipartite nonlocality
Most Bell-like inequalities proposed in the context of continuous variables hinge on a discretization process [57, 58, 59, 60, 61, 62]. Indeed, since the set of outcomes for a given observable is typically unbounded in continuous variable systems, it seems a priori difficult to derive an upper bound of the expectation value of a Bell-type observable. It is not impossible, though: relying on the Fine–Abramsky–Brandenburger theorem [63, 64], the authors of Refs. [65, 66] were able to derive continuous Bell inequalities for continuous and unbounded observables. However, to our knowledge, no practical (theoretical or experimental) use of such fully continuous approaches has been successfully implement so far. A method to circumvent this difficulty is based on discretization and consists of using a dichotomic binning of the outcome results. In this case, by using observables which rely on continuous measurements but can only take a finite number of outcomes, one can construct a Bell inequality similar to those derived for discrete variables. A well-known example of such observables are the so-called pseudo-spin operators: they live in an infinite-dimensional Hilbert space but the outcome of their measurement is either 1-1 or +1+1.
@@
[…]
@@
[…]
where $\boldsymbol{a}$, $\boldsymbol{a}^{\prime}$, $\boldsymbol{b}$ and $\boldsymbol{b}^{\prime}$ are unit vectors. Given a unit vector $\boldsymbol{n}$ it is easily checked that $(\boldsymbol{n}\cdot\hat{\boldsymbol{\Pi}}{}^{(j)})^{2}=\mathbb{1}$, meaning that the Hermitian operator $\boldsymbol{n}\cdot\hat{\boldsymbol{\Pi}}{}^{(j)}$ has eigenvalues $\pm 1$. It then follows from direct inspection that, if one believes in local realism, one should expect that a measure of the operator $\hat{\mathscr{B}}{}^{(i|j)}$ yields a result $\pm 2$ [73]. The tenant of local realism is thus violated when $\langle\hat{\mathscr{B}}{}^{(i|j)}\rangle>2$, whereas Cirel’son bound [74] imposes $\langle\hat{\mathscr{B}}{}^{(i|j)}\rangle\leq 2\sqrt{2}$. Since the modes 0 and 1 are not entangled the quantity $\langle\hat{\mathscr{B}}{}^{(0|1)}\rangle$ is always lower than 2 and its computation is of no interest to us. For attempting to violate as much as possible Bell inequality one should consider the modes 0 and 2 (or 1 and 2) and look for an arrangement of the four measurement directions $\boldsymbol{a}$, $\boldsymbol{a}^{\prime}$, $\boldsymbol{b}$ and $\boldsymbol{b}^{\prime}$ which maximizes $\langle\hat{\mathscr{B}}{}^{(i|2)}\rangle$ ($i=0$ or 1). This procedure is explained in Appendix F and makes it possible to analytically compute the quantity
@@
Figure 3: Plot of $B^{(i|2)}$ (solid lines) and $\Lambda^{(i|2)}$ (dashed lines) as functions of $\omega$ for the waterfall configuration with $m_{d}=2.9$ and different temperatures. We only consider the range of frequencies $\omega<\Omega$ for which the vacua of the outgoing and ingoing modes do not coincide. The value of the temperature is indicated in units of $g\,n_{u}=mc_{u}^{2}$. Upper plot: $i=0$, lower plot $i=1$. Non separability of modes $i$ and $2$ is achieved when $\Lambda^{(i|2)}>0$. Bell inequality is violated when $B^{(i|2)}>2$. The inset in the upper plot is a blow-up of the region $1.8\leq B^{(0|2)}\leq 2.1$ and $0.1\leq\omega/\Omega\leq 1$ showing that the reduced state $(0|2)$ does not violate Bell inequality at temperatures $T=0.1$ and 0.2.
@@
Fig. 3 indicates that, in the waterfall configuration we consider (with $m_{d}=2.9$), the entanglement is lower and the violation of Bell inequality less significant for the correlations among modes 1 and 2 than for the correlations among modes 0 and 2. However, this is not always the case. This point is illustrated in Fig. 4 which displays, for all waterfall configurations, the zero temperature values of ${\rm max}_{\omega}B^{(i|2)}$ and ${\rm max}_{\omega}\Lambda^{(i|2)}$ (for $i=0$ and 1) as functions of the upper Mach number $m_{u}$ (the maximisation is performed at fixed $m_{u}$, for $\omega\in[0,\Omega]$). All the possible waterfall configurations are considered since $m_{u}$ spans the whole interval $[0,1]$. In this figure we aim at evaluating the largest amount of entanglement and nonlocality reached in each configuration. This is the reason why we plot the maximum values taken by the quantities $\Lambda^{(i|2)}$ and $B^{(i|2)}$ over the energy interval $0,\Omega] [$where $\Omega$ is defined by Relation (7)] since this is only for energies in this interval that spontaneous emission of quasi-particles occurs.
Figure 4: Zero temperature value of the CHSH parameter and of the PPT measure characterizing non separability of modes $i$ (=0=0 and 1) and 2 for the waterfall configuration. The maximal value reached by these quantities over the interval $\omega\in[0,\Omega]$ is plotted as a function of the upper Mach number $m_{u}$ which (as explained in Appendix A) characterizes a given configuration. The values of $B^{(1|2)}$ for $m_{u}\leq 0.01$ are not indicated because of lack of numerical precision. The upper bounds of $B^{(i|2)}$ and $\Lambda^{(i|2)}$ ($\sqrt{8}$ and 1, respectively) are indicated with filled dots.
@@
Fig. 4 shows that for values of $m_{u}$ larger than 0.6, the entanglement is mainly concentrated between modes 2 and 0, i.e., between the Hawking quantum and the partner. This is indicated by the fact that both the PPT measure and the CHSH parameter significantly point to nonseparability and nonlocality between these two modes. For $m_{u}\leq 0.2$ instead, the figure shows that entanglement is concentrated between the partner and the companion (modes 2 and 1, respectively). A plot similar to the one of Fig. 4, but where the quantities are evaluated at finite temperature, enables to evaluate the resilience of entanglement and non-separability to an increase of temperature. This check is performed in Fig. 5 which shows that, whereas at $T=0.2\,gn_{u}$ the PPT measure is not dramatically affected, the CHSH parameters $B^{(0|2)}$ and $B^{(1|2)}$ no longer show evidences of violation of Bell inequality, except in the $(1|2)$ sector for waterfall configurations with $m_{u}\lesssim 0.15$ and, in a lesser extent, in the $(0|2)$ sector for $m_{u}\gtrsim 0.85$.
@@
where $\hat{\mathscr{B}}^{\prime}{}^{(j|k)}(\omega)$ is the same as $\hat{\mathscr{B}}{}^{(j|k)}(\omega)$ defined in (18) with the primes reversed, and $\boldsymbol{c}$ and $\boldsymbol{c}^{\prime}$ are normalized vectors (as well as $\boldsymbol{a}$, $\boldsymbol{a}^{\prime}$, $\boldsymbol{b}$, and $\boldsymbol{b}^{\prime}$ involved in the definition of $\hat{\mathscr{B}}{}^{(j|k)}$). Expanding expression (20) shows that $\hat{\mathscr{S}}{}^{(i|j|k)}$ is invariant upon a permutation of its indices, provided the names of the unit vectors $(\boldsymbol{a},\boldsymbol{b},\boldsymbol{c})$ and $(\boldsymbol{a}^{\prime},\boldsymbol{b}^{\prime},\boldsymbol{c}^{\prime})$ undergo the same permutation. In the following we arbitrarily chose the order $(i,j,k)=(0,1,2)$.
Similarly to what occurs for the two-mode operator, the principle of local realism, if correct, should predict that a measure of the operator $\hat{\mathscr{S}}{}^{(0|1|2)}$ yields results $\pm 2$. The system violates this principle when the average of the operator (20) is larger than two. This is often referred to as violation of Svetlichny inequality. It is important to note that the observable $\hat{\mathscr{S}}{}^{(0|1|2)}$ is specially designed in order to be sensitive to genuine tripartite nonlocality (see discussions in Refs. [77, 79, 78, 23] and references therein): a tripartite system can involve nonlocal correlations between any of its bipartitions and still not violate the Svetlichny inequality $\langle\hat{\mathscr{S}}{}^{(0|1|2)}\rangle<2$. A simple example of a system which displays two mode nonlocality but does not pass the Svetlichny test is presented in Appendix C [see Eq. (65) and the discussion below].
@@
Figure 7: Same as Fig. 6 for a temperature $T=0.05\,gn_{u}$. At variance with the zero temperature situation the tripartite measure (21) is here always less than or equal to 2.
@@
All the other expectation values of products of three components of the $\hat{\boldsymbol{\Pi}}$ operator are zero. Since all the operators in (22) have only $\pm 1$ as eigenvalues, this means that each term must reach its extremal value and the vacuum mode of zero frequency $|0_{\omega=0}\rangle^{\rm in}$ must therefore be an eigenstate of the operators appearing in (22):
@@
Relations (23) define a state which exhibits the GHZ paradox, contradicting local hidden variable theories by means of a single measurement; cf. e.g., the discussion in [80]. In the same line, the above noted fact that our system saturates the upper bound $2\sqrt{2}$ of the Svetlichny parameter is a characteristic also shared by GHZ states. Another resemblance lies in the fact that the GHZ states possess maximal tripartite entanglement, quantified by the residual tangle [81, 82] while similarly, in the domain of temperature and energy where Eqs. (23) hold, the residual contangle [83] which measures genuine tripartite entanglement in our continuous variable system diverges [22]. Despite these clear similarities, there is a significant difference between the state $|0_{\omega=0}\rangle^{\rm in}$ and the GHZ states commonly considered in quantum information theory: they have different behaviors upon partial tracing. As is well known, taking the partial trace over one of the three modes of a GHZ state yields an unentangled mixed state. At variance, partial tracing the state $|0_{\omega=0}\rangle^{\rm in}$ over modes 0 or 1 leads to an entangled two-mode mixed state, as discussed in Sec. III (see, e.g., the values $\Lambda^{(0|2)}(\omega=0)$ and $\Lambda^{(1|2)}(\omega=0)$ in Fig. 3). As we now argue, the explanation of this conundrum lies in the fact that, whereas the GHZ states are usually build on qubits, the state we consider is an infinite sum of degenerate GHZ states of a continuous variable system.
To tackle this issue it is convenient to expand the state $|0_{\omega=0}\rangle^{\rm in}$ over the eigenstates of the operators $\hat{\Pi}_{x}^{(j)}$ ($j=0$, 1 or 2). As discussed in Appendix D these eigenstates can be written as $|x^{\pm}_{n}\rangle_{j}$: they are labeled by their eigenvalue ($\pm 1$) plus another integer index ($n$ in the above expression) associated to the infinite degeneracy of both eigenvalues. Indeed, at variance with what occurs for a regular spin operator, the projection of the pseudo-spin (16) over a given axis (here $\hat{\Pi}_{x}$) has infinitely degenerate eigenvalues, i.e., there exists an infinite number of mutually orthogonal eigenstates with the same eigenvalue (+1+1 or 1-1):
@@
[…]